# The False Promise of Zero-Shot Super- Resolution in Machine-Learned Operators

### v1

@@ Abstract
A core challenge in scientific machine learning, and scientific computing more generally, is modeling continuous phenomena which (in practice) are represented discretely. Machine-learned operators (MLOs) have been introduced as a means to achieve this modeling goal, as this class of architecture can perform inference at arbitrary resolution. In this work, we evaluate whether this architectural innovation is sufficient to perform “zero-shot super-resolution,” namely to enable a model to serve inference on higher-resolution data than that on which it was originally trained. We comprehensively evaluate both zero-shot sub-resolution and super-resolution (i.e., multi-resolution) inference in MLOs. We decouple multi-resolution inference into two key behaviors: 1) extrapolation to varying frequency information; and 2) interpolating across varying resolutions. We empirically demonstrate that MLOs fail to do both of these tasks in a zero-shot manner. Consequently, we find MLOs are not able to perform accurate inference at resolutions different from those on which they were trained, and instead they are brittle and susceptible to aliasing. To address these failure modes, we propose a simple, computationally-efficient, and data-driven multi-resolution training protocol that overcomes aliasing and that provides robust multi-resolution generalization.
@@ 1 Introduction
where $M$ is an approximation of the PDE’s solution operator, $S_{1}$ is the input state of the system, and $S_{2}$ is the predicted state. Central to this problem formulation is that continuous physical systems must be sampled and, therefore, modeled discretely. For a discrete model, $M$, to be useful in representing phenomena of different scales, scientists require the ability to use it at different resolutions accurately. For example, when modeling fluid flow, scientists often use adaptive mesh refinement (Berger & Oliger, 1984), a technique that increases simulation resolution in areas that require high accuracy (e.g., regions of turbulence), and coarsens the resolution in less critical regions.
Traditionally, the approximation $M$ is constructed by numerical methods which, by design, can be employed at arbitrary discretization (Forrester et al., 2008; Cozad et al., 2014; Asher et al., 2015; Sudret et al., 2017; Alizadeh et al., 2020; Kudela & Matousek, 2022). However, numerical methods are computationally expensive. Alternatively, machine-learned operators (MLOs), a class of data-driven machine learning (ML) models which parameterize the solution operator to families of PDEs, have been proposed (Raissi et al., 2019; Li et al., 2020a; Lu et al., 2021; Kovachki et al., 2023; Raonic et al., 2023). Although querying MLOs at arbitrary discretization is computationally inexpensive, it is not obvious that this can be done accurately. The Fourier Neural Operator (FNO) (Li et al., 2020a), a specific MLO, claimed to address the discretization challenge in a zero-shot manner (Li et al., 2020a; Tran et al., 2021; George et al., 2024; Li et al., 2024b; Azizzadenesheli et al., 2024). The claim is that an FNO can be trained at resolution $m$ and can then serve accurate inference at resolution $n>m$, without training on additional high resolution data. In machine learning terminology, this is zero-shot super-resolution. This claim of zero-shot super-resolution, if true, would be especially attractive in settings where generating, and training on, high-resolution data is computationally expensive.
@@
In this paper, we evaluate the claim of zero-shot super- (and sub-)resolution inference in MLOs. We document a substantial disparity in model performance across data of different discretizations, suggesting that MLOs are generally incapable of accurate inference at resolutions greater than, or less than, their training resolution (i.e., zero-shot multi-resolution inference). Instead, we find that MLOs often misrepresent unseen frequencies and incorrectly infer their behavior in data whose discretization differs from its training discretization, i.e., they exhibit a form of aliasing (see Fig. 1). In addition, we study two previously proposed solutions: (i) physics-informed optimization constraints (Li et al., 2024b); and (ii) band-limited learning (Raonic et al., 2023; Gao et al., 2025). We find that neither enables zero-shot multi-resolution, as they do not address the central issue: MLOs, like all machine-learned models, cannot typically generalize beyond their training data (Yang et al., 2023; Liu et al., 2020; Krueger et al., 2021); and changing the discretization of test data constitutes a (specific) form of out-of-distribution inference. We establish that the discretization at which MLOs are trained impacts the discretization at which they accurately model the system.
@@
    We assess the ability of trained MLOs to generalize beyond their training resolution. We demonstrate that MLOs struggle to perform accurate inference at resolutions higher than or lower than the resolution on which they were trained, and instead exhibit a form of aliasing. Based on these results, we conclude that accurate zero-shot multi-resolution inference is brittle and unreliable. (See Sec. 3.)
@@
    We evaluate two intuitive approaches—incorporating physics-informed constraints during training, as well as performing band-limited learning—and we find that neither approach enables reliable multi-resolution generalization. (See Sec. 4.)
@@
    We propose and test multi-resolution training, where we include training data of varying resolutions (in particular, a small amount of expensive higher-resolution data and a larger amount of cheaper lower-resolution data), and we show that multi-resolution inference improves substantially, without a significant increase in training cost. (See Sec. 5.)
@@ 2 Background on Signal Processing and Aliasing
Figure 2: Accurate multi-resolution inference requires both interpolation and extrapolation. Original: Signal is sampled at a rate greater than its Nyquist frequency. Interpolation: Adapting to new sampling rates of a given signal. Extrapolation: Adapting to new frequency information under constant sampling rate. Super-Resolution: Sampling a system at a higher rate, which enables the capture of higher frequency information (interpolation and extrapolation). Aliasing: High-frequency information is misrepresented as a low-frequency information due to insufficient sampling.
We start by discussing the practice of training ML models to represent continuous systems via discrete data. Then, we outline the implications of aliasing in ML as it relates to multi-resolution inference. Finally, we formally define “zero shot multi-resolution” inference in a discrete context.
Discrete Representations of Continuous Systems. The fundamental challenge in discretely representing continuous systems lies in the choice of sampling rate. The Whittaker–Nyquist–Shannon sampling theorem established that, given a sampling rate $r$, the largest resolved frequency is $r/2$ (Unser, 2002; Shannon, 1949; Whittaker, 1928). Thus, ML models will be trained on discrete representations, where only some of the frequencies are fully resolved. Resolving higher-order frequencies greater than $r/2$, consequently, becomes an out-of-distribution task. Aligning these discrete models’ predictions with the underlying continuous system is a non-trivial open problem (Krishnapriyan et al., 2021; Queiruga et al., 2020; Ott et al., 2021; Ren et al., 2023; Takamoto et al., 2022; Subel et al., 2022; Chattopadhyay et al., 2024).
Aliasing. When sampling a continuous signal at rate $r$, aliasing occurs when frequency components greater than $r/2$ are projected onto lower frequency basis functions (see Fig. 2) (Gruver et al., 2022). When this happens, content with frequency $n>r/2$ is observed as a lower frequency contribution:
@@
Within an ML context, when inferring at different discretizations of a given signal, aliasing can manifest as the divergence between the energy spectrum of the model prediction and the expected output. Aliasing indicates a model’s failure to fit the underlying continuous system.
Zero-shot multi-resolution inference. We define multi-resolution inference as the ability to do inference at multiple resolutions (e.g., sub-resolution and super-resolution). The zero-shot multi-resolution task employs an ML model, which is trained on data with some resolution, and then tested on data with a different resolution. Zero-shot multi-resolution inference raises two important questions with respect to the generalization abilities of trained models (see Fig. 2):
@@
    Resolution Interpolation. How do models behave when the frequency information in the data remains fixed, but its sampling rate changes from training to inference? Can the model interpolate the fully-resolved signal to varying resolutions?
@@ 3 Assessing Multi-Resolution Generalization
In this section, we examine the zero-shot multi-resolution abilities of FNO, an architecture for which this claim has previously been made (Li et al., 2020a). We study the multi-resolution inference task from an out-of-distribution data perspective by decoupling what it means for a model to perform inference at a resolution different from that used during training. Specifically, in Sec. 3.1, we assess whether models trained on a system sampled at rate $r$ are capable of both interpolating accurately to new sampling rates (Fig. 2($b$)) and extrapolating accurately to additional/fewer frequencies present in data (Fig. 2($c$)). We systematically test an FNO’s ability to do both objectives and show neither are achieved. In light of these failure modes, in Sec. 3.2, we then assess the spatial zero-shot sub- and super-resolution capabilities of FNOs, and we show the claim does not hold (Fig. 2($d$)).
We evaluate FNO on three standard scientific datasets: Darcy, Burgers, and Navier Stokes described in Appendix A. For each dataset we optimize FNO hyperparameters via grid-search as described in Appendix B.
Figure 3: Resolution Interpolation. Four FNOs are trained on Darcy data at resolutions {16, 32, 64, 128} from left to right with constant frequency information (low-pass limit of 8ff). We test whether each model can generalize to data with varying resolution, visualizing the normalized spectra of the residuals averaged across test data. Notice that the residual spectra (error) increases substantially in the higher frequencies. (Lower residual energy at all frequencies is better.)
@@ 3.1 Breaking Down Multi-Resolution Capabilities
Resolution Interpolation. We assess whether an FNO trained on data of a specific resolution can generalize to data of both lower and higher resolution, under fixed frequency information. Specifically, we keep the set of populated frequencies constant in the train and test data, while varying the resolution of the test data. We do this by applying a low-pass filter to all data at the highest resolution, and then down-sampling as needed. The sampling rates of all data are large enough to resolve all remaining frequencies.
@@
We begin with a simple experiment: we train an FNO on a Darcy flow dataset at resolution $N=16$ and assess the trained model’s performance across test datasets at varying resolutions $\{16,32,64,128\}$, all low-pass filtered with limit 8ff, where $f$ is the frequency unit $2\pi/N$. In Fig. 3($a$), we visualize the average spectral energy of the model predictions, normalized by the spectral energy of the unfiltered ground truth for each test dataset. For the test data with resolutions that are different from the training data, we observe sharp increases in their residual’s energy spectrum in frequencies greater than 8ff. This is especially concerning since the model was never trained on nor shown inference data containing frequencies greater than 8ff. In other words, FNOs, trained in a zero-shot manner, fail to reliably interpolate to varying resolutions.
@@
In Figs. 3($b$-$d$), we repeat the same experiment at varying training resolutions (e.g, 32, 64, 128), with low-pass limit 8ff. We notice that at each training resolution, the model consistently and incorrectly assigns high energy in frequencies greater than 8ff across all test resolutions. We perform corresponding experiments for the Burgers dataset with low-pass limit 64ff and resolutions $\in$ {128, 256, 512, 1024} and Navier Stokes dataset with low-pass limit 32ff and resolutions $\in$ {64, 128, 255, 510} and observe the same failure mode (see Appendix C: Figs. 11-12). We conclude that changing resolution at test time is akin to out-of-distribution inference: the model is never trained on data with a broad range of sampling rates, and consequently, it fails to generalize.
@@
Figure 4: Information Extrapolation. Four FNOs trained on Darcy data of resolution 128 (constant sampling rate) and low-pass filtered with limits {8, 16, 32, 64}$f$ (varying frequency information) from left to right. We test whether each model can generalize to data with varying frequency information, visualizing the normalized spectra of the residuals averaged across test data. Notice that the residual spectra (error) increases substantially in the higher frequencies. (Lower residual energy at all frequencies is better.)
@@
Information Extrapolation. We assess whether an FNO trained on data containing a fixed set of frequencies can generalize to data containing both fewer and additional frequencies, under fixed resolution. Specifically, we keep the resolution constant but vary the number of populated frequencies in the train and test datasets by applying varying low-pass filters to data at a fixed resolution.
@@
In Figs. 4($b$-$d$), we repeat the same experiment at varying training data low-pass filter limits (e.g, 16, 32, 64ff). Each model consistently and incorrectly assigns high energy in the high frequencies, regardless of whether the test data contained any high-frequency information. This is a concerning failure mode indicating FNOs do not generalize both in the presence of frequencies greater than, and the absence of frequencies less than, those which were present in their training data. We perform corresponding experiments for the Burgers dataset with resolution 1024 and low-pass limits $\in$ {64, 128, 256, 512}$f$ and Navier Stokes dataset with resolution 510 and low-pass limits $\in$ {8, 16, 32, 255}$f$, and observe the same failure mode (see Appendix C: Figs. 13-14). We conclude that varying the frequency information at test time is effectively out-of-distribution inference as the model was not trained on data with such variation in frequency information, and, therefore, it fails to generalize.
@@ 4 Evaluating Potential Corrective Methods
In this section, we assess two previously proposed strategies for accurate zero-shot multi-resolution inference: physics-informed optimization objectives (Li et al., 2024b) and band-limited learning (Raonic et al., 2023; Gao et al., 2025). We find that neither enables accurate multi-resolution inference.

### v2

@@ Abstract
A core challenge in scientific machine learning, and scientific computing more generally, is modeling continuous phenomena which (in practice) are represented discretely. Machine-learned operators (MLO) have been introduced as a means to achieve this modeling goal, as this class of architecture can perform inference at arbitrary resolution. In this work, we evaluate whether this architectural innovation is sufficient to perform “zero-shot super-resolution,” namely to enable a model to serve inference on higher-resolution data than that on which it was originally trained. We comprehensively evaluate both zero-shot sub-resolution and super-resolution (i.e., multi-resolution) inference in MLOs. We decouple multi-resolution inference into two key behaviors: 1) extrapolation to varying frequency information; and 2) interpolating across varying resolutions. We empirically demonstrate that MLOs fail to do both of these tasks in a zero-shot manner. Consequently, we find MLOs are not able to perform accurate inference at resolutions different from those on which they were trained, and instead they are brittle and susceptible to aliasing. To address these failure modes, we propose a simple, computationally-efficient, and data-driven multi-resolution training protocol that overcomes aliasing and that provides robust multi-resolution generalization.
@@ 1 Introduction
where $M$ is an approximation of the PDE’s solution operator, $S_{1}$ is the input state of the system, and $S_{2}$ is the predicted state. Central to this problem formulation is that continuous physical systems must be sampled and, therefore, modeled discretely. For a discrete model, $M$, to be useful in representing phenomena of different scales, scientists require the ability to use it at different resolutions accurately. For example, when modeling fluid flow, scientists often use adaptive mesh refinement (Berger and Oliger, 1984), a technique that increases simulation resolution in areas that require high accuracy (e.g., regions of turbulence), and coarsens it in less critical regions.
Traditionally, the approximation $M$ is constructed by numerical methods which, by design, can be employed at arbitrary discretization (Forrester et al., 2008; Cozad et al., 2014; Asher et al., 2015; Sudret et al., 2017; Alizadeh et al., 2020; Kudela and Matousek, 2022). However, numerical methods are computationally expensive. Alternatively, machine-learned operators (MLOs), a class of data-driven machine learning (ML) models which parameterize the solution operator to families of PDEs, have been proposed (Raissi et al., 2019; Li et al., 2020a; Lu et al., 2021b; Kovachki et al., 2023; Raonic et al., 2023). Although querying MLOs at arbitrary discretization is computationally inexpensive, it is not obvious that this can be done accurately. The Fourier Neural Operator (FNO) (Li et al., 2020a), a specific MLO, claimed to address the discretization challenge in a zero-shot manner (Li et al., 2020a; Tran et al., 2021; George et al., 2024; Li et al., 2024b; Azizzadenesheli et al., 2024). The claim is that FNO can be trained at resolution $m$ and then serve accurate inference at resolution $n>m$, without training on additional high resolution data e.g., zero-shot super-resolution. This claim of zero-shot super-resolution, if true, is especially attractive in settings where generating, and training on, high-resolution data is computationally expensive.
@@
In this paper, we evaluate the claim of zero-shot super- (and sub-)resolution inference in MLOs. We document a substantial disparity in model performance across data of different discretizations, suggesting that MLOs are generally incapable of accurate inference at resolutions greater or less than their training resolution (i.e., zero-shot multi-resolution inference). Instead, we find that MLOs often misrepresent unseen frequencies and incorrectly infer their behavior in data whose discretization differs from its training discretization, i.e., they exhibit a form of aliasing (Fig. 1). In addition, we study two previously proposed solutions: (i) physics-informed optimization constraints (Li et al., 2024b) and (ii) band-limited learning (Raonic et al., 2023; Gao et al., 2025). We find that neither enables zero-shot multi-resolution, as they do not address the central issue: MLOs, like all machine-learned models, cannot typically generalize beyond their training data (Yang et al., 2023; Liu et al., 2020; Krueger et al., 2021). We establish that the discretization at which MLOs are trained impacts the discretization at which they accurately model the system.
@@
    We assess the ability of trained MLOs to generalize beyond their training resolution. We demonstrate that MLOs struggle to perform accurate inference at resolutions higher or lower than which are they trained on, and instead they exhibit aliasing. Based on these results, we conclude that accurate zero-shot multi-resolution inference is unreliable (Sec. 3).
@@
    We evaluate two intuitive approaches—incorporating physics-informed constraints during training, as well as performing band-limited learning—and we find that neither approach enables reliable multi-resolution generalization. (Sec. 4).
@@
    We propose and test multi-resolution training, where we include training data of varying resolutions (in particular, a small amount of expensive higher-resolution data and a larger amount of cheaper lower-resolution data), and we show that multi-resolution inference improves substantially, without a significant increase in training cost (Sec. 5).
@@ 2 Background on Signal Processing and Aliasing
Figure 2: Accurate multi-resolution inference requires both interpolation & extrapolation. Original: Signal is sampled at a rate greater than its Nyquist frequency. Interpolation: Adapting to new sampling rates of a given signal. Extrapolation: Adapting to new frequency information under constant sampling rate. Super-Resolution: Sampling a system at a higher rate which enables the capture of higher frequency information (interpolation & extrapolation). Aliasing: High-frequency information is misrepresented as a low-frequency information due to insufficient sampling.
We start by discussing the practice of training ML models to represent continuous systems via discrete data. Next, we outline the implications of aliasing in ML as it relates to multi-resolution inference. Finally, we formally define “zero shot multi-resolution” inference in a discrete context.
Discrete Representations of Continuous Systems. The fundamental challenge in discretely representing continuous systems lies in the choice of sampling rate. The Whittaker–Nyquist–Shannon sampling theorem established that given a sampling rate $r$, the largest resolved frequency is $r/2$ (Unser, 2002; Shannon, 1949; Whittaker, 1928). Thus, ML models will be trained on discrete representations where only some of the frequencies are fully resolved. Resolving higher-order frequencies greater than $r/2$, consequently, becomes an out-of-distribution task. Aligning these discrete models’ predictions with the underlying continuous system is an open problem (Krishnapriyan et al., 2021; Queiruga et al., 2020; Ott et al., 2021; Ren et al., 2023; Takamoto et al., 2022; Subel et al., 2022; Chattopadhyay et al., 2024).
Aliasing. When sampling a continuous signal at rate $r$, aliasing occurs when frequency components greater than $r/2$ are projected onto lower frequency basis functions (Fig. 2) (Gruver et al., 2022). Thus, content with frequency $n>r/2$ is observed as a lower frequency contribution:
@@
In an ML context, when inferring at different discretizations of a given signal, aliasing can manifest as the divergence between the energy spectrum of the model prediction and the expected output. Aliasing indicates a model’s failure to fit the underlying continuous system.
Zero-shot multi-resolution inference. We define multi-resolution inference as the ability to do inference at multiple resolutions (e.g., sub- and super-resolution). The zero-shot multi-resolution task employs an ML model, which is trained on data with some resolution and tested on data with a different resolution. Zero-shot multi-resolution inference raises two important questions with respect to the generalization abilities of trained models (see Fig. 2):
@@
    Resolution Interpolation. How do models behave when the frequency information in the data remains fixed, but its sampling rate changes from training to inference? Can the model interpolate the fully resolved signal to varying resolutions?
@@ 3 Assessing multi-resolution Generalization
We examine the zero-shot multi-resolution abilities of FNO, an architecture for which this claim has been previously made (Li et al., 2020a). We study the multi-resolution inference task from an out-of-distribution data perspective by decoupling what it means for a model to perform inference at a resolution different from that used during training. Specifically, in Sec. 3.1, we assess whether models trained on a system sampled at rate $r$ are capable of both interpolating accurately to new sampling rates (Fig. 2($b$)) and extrapolating accurately to additional/fewer frequencies present in data (Fig. 2($c$)). We systematically test an FNO’s ability to do both objectives and show neither are achieved. In light of these failure modes, in Sec. 3.2, we then assess the spatial zero-shot sub- and super-resolution capabilities of FNOs and show the claim does not hold (Fig. 2($d$)).
We evaluate FNO on three standard scientific datasets: Darcy, Burgers, and Navier Stokes. For each dataset we optimize FNO hyperparameters via grid-search as described in Appendix B.
Figure 3: Resolution Interpolation. Four FNOs are trained on Darcy data at resolutions {16, 32, 64, 128} from left to right with constant frequency information (low-pass limit of 8ff), and are tested on varying resolutions. We assess if each model can generalize to data with varying sampling rate. We visualize spectra of the normalized residuals across test data. Notice, residual spectra (error) increases substantially in the low frequencies. Lower residual energy at all frequencies is better.
@@ 3.1 Breaking Down Multi-Resolution Capabilities
Resolution Interpolation. We assess if an FNO trained on data of a specific resolution can generalize to data of both lower and higher resolution under fixed frequency information. Specifically, we keep the set of populated frequencies constant in the train and test data, while varying the resolution of the test data. We do this by applying a low-pass filter to all data at the highest resolution, and then down-sampling as needed. The sampling rates of all data are large enough to resolve all remaining frequencies.
@@
We begin with a simple experiment: we train an FNO on a Darcy flow dataset at resolution $N=16$ and assess the trained model’s performance across test datasets at varying resolutions $\{16,32,64,128\}$, all low-pass filtered with limit 8ff where $f$ is the frequency unit $2\pi/N$. In Fig. 3($a$), we visualize the average spectral energy of the model predictions normalized by the spectral energy of the unfiltered ground truth for each test dataset. For the test data with resolutions that are different from the training data, we observe sharp increases in their residual’s energy spectrum in frequencies greater than 8ff. This is especially concerning since the model was never trained on nor shown inference data containing frequencies greater than 8ff. In other words, FNOs, trained in a zero-shot manner, fail to reliably interpolate to varying resolutions.
@@
In Figs. 3($b$-$d$), we repeat the same experiment at varying training resolutions (e.g, 32, 64, 128) with low-pass limit 8ff. We notice that at each training resolution, the model consistently and incorrectly assigns high energy in frequencies greater than 8ff across all test resolutions. We perform corresponding experiments for the Burgers dataset with low-pass limit 64ff and resolutions $\in$ {128, 256, 512, 1024} and Navier Stokes dataset with low-pass limit 32ff and resolutions $\in$ {64, 128, 255, 510} and observe the same failure mode (Appendix C: Figs. 11-12). We conclude that changing resolution at test time is akin to out-of-distribution inference: the model is never trained on data with a broad range of sampling rates and consequently fails to generalize.
@@
Figure 4: Information Extrapolation. Four FNOs trained on Darcy data of resolution 128 (constant sampling rate) and low-pass filtered with limits {8, 16, 32, 64}$f$ (varying frequency information) from left to right. Test if each model can generalize to data with varying frequency information. Visualizing spectrum of the normalized residuals across test data. Notice, residual spectra (error) increases substantially in the high frequencies. Lower residual energy at all frequencies is better.
@@
Information Extrapolation. We assess if an FNO trained on data containing a fixed set of frequencies can generalize to data containing both fewer and additional frequencies under fixed resolution. Specifically, we keep the resolution constant but vary the number of populated frequencies in the train and test datasets by applying varying low-pass filters to data at a fixed resolution.
@@
In Figs. 4($b$-$d$), we repeat the same experiment at varying training data low-pass filter limits (e.g, 16, 32, 64ff). Each model consistently and incorrectly assigns high energy in the high frequencies regardless of whether the test data contained any high-frequency information. This is a concerning failure mode indicating FNOs do not generalize both in the presence of frequencies greater than, and the absence of frequencies less than, what was present in their training data. We perform corresponding experiments for the Burgers dataset with resolution 1024 and low-pass limits $\in$ {64, 128, 256, 512}$f$ and Navier Stokes dataset with resolution 510 and low-pass limits $\in$ {8, 16, 32, 255}$f$ and observe the same failure mode (Appendix C: Figs. 13-14). We conclude that varying the frequency information at test time is effectively out-of-distribution inference as the model was not trained on data with such variation in frequency information and, therefore, fails to generalize.
@@ 4 Evaluating Potential Corrective Methods
Here we assess two previously proposed strategies for accurate zero-shot multi-resolution inference: physics-informed optimization objectives (Li et al., 2024b) and band-limited learning (Raonic et al., 2023; Gao et al., 2025). We find that neither enables accurate multi-resolution inference.